Farris Industrial purchased a machine five years ago at a cost of $164,900. The machine is being depreciated using the straight-line method over eight years. The tax rate is 21 percent and the discount rate is 14 percent. If the machine is sold today for $42,500, what will the aftertax salvage value be?

A.

$31,794.72

B.

$49,268.13

C.

$38,439.13

D.

$46,560.88

Kustom Cars purchased a fixed asset two years ago for $39,000 and sold it today for $19,000. The assets are classified as 5-year property for MACRS. The MACRS table values are 0.2000, 0.3200, 0.1920, 0.1152, 0.1152, and 0.0576 for Years 1 to 6, respectively. What is the net cash flow from the salvage value if the tax rate is 21 percent?

$18,941

$18,020

$18,902

$19,000
Answer #1
As per chegg guidelines we answer one question per post. Kindly post remaining questions in next post
D.
$46,560.88

Dear Student
Thank you for using Chegg
Please find below the answer
Statementshowing Computations
PaticularsAmount
Cost of Machine        164,900.00
Life in years                     8.00
Depreciation per year= 164900/8           20,612.50
Depreciation for 5 Years = 20,612.50*5        103,062.50
Carrying Value = 164900 - 103,062.50           61,837.50
Sale Value           42,500.00
Loss on Sale = 61,837.50 - 42,500           19,337.50
Tax savings on loss = 19337.50 *21%             4,060.88
aftertax salvage value = 42500 + 4060.88           46,560.88


answered by: ANURANJAN SARSAM